OpenStudy (anonymous):

let f(x)=(6x-7)^5 find f"(x)

OpenStudy (hba):

differentiate it 2 times

OpenStudy (anonymous):

is that f'(x) as in the derivative? f'(x)=df/dx=5(6x-7)^4 * (6)

OpenStudy (callisto):

f(x)=(6x-7)^5 f'(x)=5(6)(6x-7)^4 = 30(6x-7)^4 f''(x)= diff. f'(x)

OpenStudy (hba):

he says he ahs problem with the last step only he has reached till here

OpenStudy (anonymous):

DOUBLE PRIME OH I DIDN'T SEE THAT LOL so f'(x)=30(6x-7)^4 then f''(x)=30*4(6x-7)^3=120(6x-7)^3 by chain rule again

OpenStudy (anonymous):

What will you get when you multiply the coefficient with the power @qasaali

OpenStudy (anonymous):

@qasaali if you will not interact with us then we can't help you.. What we are saying just give a reply to that..

OpenStudy (anonymous):

I said to interact here @qasaali not in the chat box..

OpenStudy (anonymous):

ow yeah!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Simply tell me what is coefficient and what is the power here after taking first derivative..

OpenStudy (anonymous):

u coefficient the is variables which are found infront of any expression, and Power is the numbers which are found at the top of varaibles or expressions as for example X^2............................... X^3 AND ETC

OpenStudy (anonymous):

You have the concept but you are not applying here that concept: See: \[f'(x) = 30(6x-7)^4\] Here tell me what is the value of coefficient and what is power here..

OpenStudy (anonymous):

Here the coefficient is 30.............................................. and the Power is the 4 up there

OpenStudy (anonymous):

water got the point

OpenStudy (anonymous):

Yes you are right.. Then multiply both of them: 30*4 = ?? Are you helping me or I am helping you?? Forget.. Just multiply them and tell me what you got..

OpenStudy (anonymous):

u r helping me

OpenStudy (anonymous):

i need an other help man i will post c it